# LEAD: Towards Learning-Based Equity-Aware Decarbonization in Ridesharing Platforms

### v1

@@ 4 The LEAD Algorithm
Figure 1: Overview of LEAD. During each batch, LEAD evaluates the long-term emissions and utilities and uses them to construct the weighting of the ILP problem of (10). After solving the optimization problem and matching ride requests with drivers, LEAD updates the value functions based on the deadhead and trip distances of the served requests in the batch.
4.1 Learning Expected Emissions and Utility
In the following, we discuss the use of online reinforcement learning to model the impact of current assignments on future assignments and evaluate the expected deadhead and trip distances of trips in different regions of the covered area. Reinforcement learning is a method where agents learn by interacting with the environment over time. At each step, the agent takes an action and receives feedback in the form of utility from the environment. This continuous interaction allows the agent to optimize its strategy based on the utility received, making it well-suited for dynamic and complex systems such as ridesharing, where decisions made at one point can influence outcomes in the future. In our work, the reinforcement learning module consists of the following elements:
@@ 4.1 Learning Expected Emissions and Utility
  • State: The state of driver $v$ during batch $b$ is represented by the driver’s location $l_{v,b}$. To facilitate data analysis and management, city space is typically discretized into a square grid system, which results in a finite number of unique states.
@@
Due to the complexity of considering multiple dependent agents within the system, we simplify the learning module by treating different agents as a single agent. In the emission-aware fair assignment module, we leverage the evaluated value functions to take appropriate actions for the different action-correlated agents. We adopt a framework that involves sequences comprising the current state, current action, reward, and next state to estimate the value functions. One effective method for learning in this scenario is Temporal Difference (TD) learning [39], which iteratively updates value functions based on the deadhead and trip distances of each completed ride request.
@@
$V_{D}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,D}|s_{0}=s\right],$ (4)
where $d_{t,D}$ is the deadhead distance of $t^{th}$ trip served by the agent, $s_{0}$ denotes the starting state of the agenet, and $\gamma\leq 1$ is the discount factor. Similarly, we define $V_{T}(s)$ as the expected trip distance a driver will travel when starting from state $s$.
$V_{T}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,T}|s_{0}=s\right],$ (5)
where $d_{t,T}$ is the trip distance of $t^{th}$ trip served by the agent.
Assume the agent is located in state $s_{t}$ and after serving the request, travels to the next state $s_{t+1}$. The TD learning update rule for the value functions based on the trip and deadhead distance of the trip are as follows:
$V_{D}(s)\leftarrow$ $V_{D}(s)+\alpha\left[d_{t,D}+\gamma V_{D}(s_{t+1})-V_{D}(s_{t})\right],$ (6)
$V_{T}(s)\leftarrow$ $V_{T}(s)+\alpha\left[d_{t,T}+\gamma V_{T}(s_{t+1})-V_{T}(s_{t})\right],$ (7)
@@
2       Calculate $E_{v,r,b}$ using Equation (4.2);
3       Calculate $U_{v,r,b}$ using Equation (4.2);
@@ 11       Update $V_{T}(s_{t})$ using Equation (7);
4.2 Batched Emission-Aware Fair Assignment
In this section, we propose the process of matching ride requests to available drivers within LEAD. LEAD performs the ride matching process in a batch setting, where each batch comprises ride requests posted after the assignments from the previous batch. Let $E_{v,r,b}$ denote the long-term expected emission produced by driver $v$ after being assigned to request $r$ within batch $b$. The value of $E_{v,r,b}$ depends on the trip and deadhead distances for serving request $r$, as well as future deadhead and trip distances traveled by driver $v$. To estimate the latter, we leverage the value function described in the previous section. Specifically, the expected emission produced by driver $v$ during and after batch $b$ can be calculated as:
$=\bigg{[}(d_{r}^{(T)}+d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})+V_{D}(q_{r}))$
$-(V_{T}(l_{v,b})+V_{D}(l_{v,b}))\bigg{]}\cdot e_{v},$ (8)
$=U_{v,r}+(d_{r}^{(T)}-d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})-V_{D}(q_{r}))$
$-(V_{T}(l_{v,b})+V_{D}(l_{v,b})).$ (9)
$\min\quad$ $\text{Emission}_{(b)}-\eta\cdot\text{Fairness}_{(b)},$ (10a)
$\text{Emission}_{(b)}:=\sum_{v\in\mathcal{V}_{b}}\sum_{r\in\mathcal{R}_{b}}E_{v,r,b}\cdot x_{v,r},$ (10b)
$\text{Fairness}_{(b)}:=-\max_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r})$
$+\min_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r}),$ (10c)
s.t., $\sum_{v\in\mathcal{V}_{b}}x_{v,r}\in\{0,1\}\quad\forall r,$ (10d)
$\sum_{r\in\mathcal{R}_{b}}x_{v,r}\in\{0,1\}\quad\forall v,$ (10e)
vars., $\quad\quad\ x_{v,r}\in\{0,1\}\quad\forall v,r,$ (10f)
The pseudocode for LEAD is outlined in Algorithm 1. Initially, LEAD evaluates the values of $E_{v,r,b}$ and $U_{v,r,b}$ for every pair of $v$ and $r$ in batch $b$ (Lines 1-3). Next, it finds the solution for the optimization problem in (10), determining which driver should serve each ride request. The final step is to update the value functions based on the actual trip and deadhead distances of the served ride requests within this batch. Specifically, for each request $r$ in the batch, LEAD updates the value functions associated with trip and deadhead distances using the update rules presented in Equations (6) and (7) (Lines 5-10). By integrating emission and fairness into a single objective, LEAD balances reducing carbon emissions with the fairness of utilities among drivers. Additionally, by considering the long-term expected emissions and utilities instead of just the current values, LEAD optimizes these metrics with a forward-looking perspective.
5 Experimental Evaluation
In this section, we conduct comprehensive experiments to evaluate the performance of LEAD using several metrics, including emissions per trip, fairness in driver’s utility, and average wait times for riders. We compare LEAD against state-of-the-art algorithms across various values for different batch durations, the fairness parameter $\eta$, and the percentages of low-emission vehicles. Below, we outline the key questions the evaluation addresses and summarize our findings.
@@ 5 Experimental Evaluation
    How does LEAD reduce emissions, especially compared to other baselines targeting emissions reduction?
    Outcome: LEAD reduces emissions by at least 73.6%, 70.5%, and 58.3% compared to CD, TORA, and LAF algorithms, respectively.
@@
    How much unfairness does LEAD incur to reduce emissions, compared to LAF that only targets fairness?
    Outcome: LEAD offers a favorable, and more importantly configurable, tradeoff between fairness and emissions reduction; it reduces emissions by 65% for a 9% drop in fairness. It also exposes the $\eta$ knob to enable navigating the tradeoff gracefully.
@@
    How does LEAD’s rider wait time performance compare to the baselines that explicitly consider reducing wait time?
@@
    Outcome: LEAD achieves at least a 40% improvement over baseline algorithms while also ensuring that improving fairness for drivers does not impact the rider’s wait time by more than 3%.
@@
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
@@
Figure 2: Emissions reduction performance: (a) emission per trip for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.1 Experimental Setup
First, we provide details on the ridesharing dataset, baseline algorithms, metrics, and the range of evaluation parameters.
Ridesharing dataset. Our experiments use a publicly available dataset from RideAustin [28], a non-profit ridesharing service based in Austin, Texas. This dataset includes approximately 1.5 million trips over ten months in 2016 and 2017, collected in Austin. It contains detailed trip information such as pick-up and drop-off coordinates of each trip, vehicle make and model, distances traveled before, during, and after each trip.
In our experiments, we use a subset of the RideAustin dataset, focusing on trips from December 1, 2016, to December 15, 2016. This subset includes 58,866 ridesharing trips served by 150 drivers. We augment this dataset by including carbon emission data for both deadhead miles and individual trips, as well as equity information for drivers and riders, using E2-RideKit, a publicly available toolkit [32]. We assume a rider cancels their request if it remains unassigned for two consecutive batches, and considered drivers available within 15 minutes in the set of available drivers of the batch. We also classified vehicles emitting under 135 g.CO2/km as low-emission (LEVs) and over 270 g.CO2/km as high-emission (HEVs). To assess the impact of replacing HEVs with LEVs, we created three dataset variants by randomly converting a certain percentage of the non-LEVs into electric vehicles (EVs), resulting in datasets with 10%, 15%, and 25% LEVs. We set the EV’s emission intensity to 63.35 g.CO2 per kilometer. [^1] We used the Tesla Model Y as a representative EV, which has an energy efficiency rating of 26 kWh/100mi. The unit distance emissions were calculated using the average carbon intensity of 408 g.CO2eq/kWh for Austin, Texas.
Comparison algorithms. We evaluate the performance of LEAD against a heuristic baseline algorithm and two state-of-the-art algorithms from prior works detailed below:
  • Closest Driver (CD): For a given batch of requests, this algorithm sequentially matches requests in the batch and greedily assigns a request to the nearest available driver to minimize the deadhead miles and the rider’s wait time.
  • LAF [35]: This algorithm has four stages: evaluating, assigning, guiding, and learning. Initially, it evaluates a bipartite graph where edge weights represent trip prices, and these weights are updated to reflect both current and future earnings. During the assigning stage, LAF utilizes a bi-objective assignment algorithm to balance the aggregate utilities of drivers and the equity in their utilities. The learning stage involves refining weight updates through reinforcement learning to guide future assignments.
    Since LAF aims to maximize total utilities and ensure equity in utilities, it uses a single value function to track the expected achievable utilities in different city regions. For adaptation to the current work, we set the utility (or price) of a ride request $r$ served by driver $v$ as the difference between the trip and deadhead distances and evaluated weights in Algorithm 1 of [35] accordingly. Additionally, during our experiments, the weight of trip utilities over different times of the day is the same.
  • TORA  [32]: TORA is an online ride-assignment algorithm that aims to balance passenger waiting times with system-wide emissions. To the best of our knowledge, it is the only existing online ride assignment algorithm that explicitly considers emissions. TORA aims to reduce passenger waiting times by initially selecting the closest available driver. It then evaluates alternative available drivers by comparing their distances and the additional deadhead emissions they would generate relative to the closest driver. For each driver, TORA calculates an “Emission-to-Distance” (E2D) ratio, which is the ratio of the difference in deadhead emissions and the deadhead miles from the two drivers to the passenger. Finally, the algorithm selects the driver with the most favorable E2D ratio. Although TORA accounts for carbon emissions during ride matching, it does not address fairness in driver utilities.
Parameter ranges. We evaluate the performance of LEAD and baseline algorithms in various experimental settings. We evaluate the impact of parameter $\eta$ and batch duration by varying $\eta$ between 0.1 g.CO2/km and 10 g.CO2/km and varying batch duration values between 2 and 15 minutes. To define unique states, we divide the city into square tiles with a width of 1 km, initialize value functions to 0, set the discount factor $\gamma=0.9$, and the learning rate $\alpha=0.025$ to align with the evaluation settings used in [35].
Performance metrics. We evaluate the performance of LEAD and comparison algorithms using three performance metrics: per trip emissions, fairness in driver’s utility, and rider waiting time. For fairness, we normalize our results to LAF outcomes when comparing different algorithms. However, in analyzing the effect of LEV penetration, we normalize the values to the LEAD without the addition of LEVs. We also report the proportion of trips assigned to low-emission vehicles (LEVs) and high-emission vehicles (HEVs), as well as the proportion of successfully matched requests.

### v2

@@ 4. The LEAD Algorithm
Figure 1. Overview of LEAD. During each batch, LEAD evaluates the long-term emissions and utilities and uses them to construct the weighting of the ILP problem of (10). After solving the optimization problem and matching ride requests with drivers, LEAD updates the value functions based on the deadhead and trip distances of the served requests in the batch.
4.1. Learning Expected Emissions and Utility
In the following, we discuss the use of online reinforcement learning to model the impact of current assignments on future assignments and evaluate the expected deadhead and trip distances of trips in different regions of the covered area. Reinforcement learning is a method where agents learn by interacting with the environment over time. At each step, the agent takes an action and receives feedback in the form of utility from the environment. This continuous interaction allows the agent to optimize its strategy based on the utility received, making it well-suited for dynamic and complex systems such as ridesharing, where decisions made at one point can influence outcomes in the future.
In this work, we focus on learning value functions to enable future-aware ride assignment. Common model-free algorithms, such as Temporal Difference (TD) learning (Sutton, 1988), Q-learning (Watkins and Dayan, 1992), and Actor-Critic methods (Konda and Tsitsiklis, 1999), can be leveraged for this task, each with its strengths and limitations. Q-learning, which estimates action-value functions, faces challenges in our case due to insufficient data for each state-action pair. Actor-Critic methods, which learn both a value function and a policy simultaneously (with the Critic estimating the value and the Actor updating the policy based on feedback), are also unsuitable as the policy is dictated by riders and cannot be altered. This coupling of policy and value functions limits their use in our context. Given these constraints, we adopt TD learning for our work. TD learning updates value estimates incrementally, making it well-suited to our data limitations. The key elements of our reinforcement learning module are as follows:
@@ 4.1. Learning Expected Emissions and Utility
  • State: The state of driver $v$ during batch $b$ is represented by the driver’s location $l_{v,b}$. To facilitate data analysis and management, city space is discretized into a square grid system, which results in a finite number of unique states.
@@
Due to the complexity of considering multiple dependent agents within the system, we simplify the learning module by treating different agents as a single agent. In the emission-aware fair assignment module, we leverage the evaluated value functions to take appropriate actions for the different action-correlated agents. We adopt a framework that involves sequences comprising the current state, current action, reward, and next state to estimate the value functions. For learning in this scenario, we use TD learning which iteratively updates value functions based on the deadhead and trip distances of each completed ride request.
@@
(4) $V_{D}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,D}|s_{0}=s\right],$
where $d_{t,D}$ is the deadhead distance of $t^{th}$ trip served by the agent, $s_{0}$ denotes the starting state of the agent, and $\gamma\leq 1$ is the discount factor. Similarly, we define $V_{T}(s)$ as the expected trip distance a driver will travel starting from state $s$.
(5) $V_{T}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,T}|s_{0}=s\right],$
where $d_{t,T}$ is the trip distance of $t^{th}$ trip served by the agent. Now, assume the agent is located in state $s_{t}$ and after serving the request, travels to the next state $s_{t+1}$. The TD learning update rule for the value functions based on the trip and deadhead distance of the trip are as follows:
(6) $V_{D}(s)\leftarrow$ $V_{D}(s)+\alpha\left[d_{t,D}+\gamma V_{D}(s_{t+1})-V_{D}(s_{t})\right],$
(7) $V_{T}(s)\leftarrow$ $V_{T}(s)+\alpha\left[d_{t,T}+\gamma V_{T}(s_{t+1})-V_{T}(s_{t})\right],$
@@
4.2. Batched Emission-Aware Fair Assignment
This section proposes matching ride requests to available drivers within LEAD. LEAD performs the ride matching process in a batch setting, where each batch comprises ride requests posted after the assignments from the previous batch. Let $E_{v,r,b}$ denote the long-term expected emission produced by driver $v$ after being assigned to request $r$ within batch $b$. The value of $E_{v,r,b}$ depends on the trip and deadhead distances for serving request $r$, as well as future deadhead and trip distances traveled by driver $v$. To estimate the latter, we leverage the value function described in the previous section. Specifically, the expected emission produced by driver $v$ during and after batch $b$ can be calculated as:
(8) $=\bigg{[}(d_{r}^{(T)}+d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})+V_{D}(q_{r}))-(V_{T}(l_{v,b})+V_{D}(l_{v,b}))\bigg{]}\cdot e_{v},$
(9) $=U_{v,r}+(d_{r}^{(T)}-d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})-V_{D}(q_{r}))-(V_{T}(l_{v,b})+V_{D}(l_{v,b})).$
(10a) $\min\quad$ $\text{Emission}_{(b)}-\eta\cdot\text{Fairness}_{(b)},$
(10b) $\text{Emission}_{(b)}:=\sum_{v\in\mathcal{V}_{b}}\sum_{r\in\mathcal{R}_{b}}E_{v,r,b}\cdot x_{v,r},$
(10c) $\text{Fairness}_{(b)}:=-\max_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r})+\min_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r}),$
(10d) s.t., $\sum_{v\in\mathcal{V}_{b}}x_{v,r}\in\{0,1\}\quad\forall r,$
(10e) $\sum_{r\in\mathcal{R}_{b}}x_{v,r}\in\{0,1\}\quad\forall v,$
(10f) vars., $\quad\quad\ x_{v,r}\in\{0,1\}\quad\forall v,r,$
The pseudocode for LEAD is outlined in Algorithm 1. Initially, LEAD evaluates the values of $E_{v,r,b}$ and $U_{v,r,b}$ for every pair of $v$ and $r$ in batch $b$ (Lines 1-3). Next, it finds the solution for the optimization problem in (10), determining which driver should serve each ride request. The final step is to update the value functions based on the actual trip and deadhead distances of the served ride requests within this batch. Specifically, for each request $r$ in the batch, LEAD updates the value functions associated with trip and deadhead distances using the update rules presented in Equations (6) and (7) (Lines 5-10). By integrating emission and fairness into a single objective, LEAD balances reducing carbon emissions with the fairness of utilities among drivers. Additionally, by considering the long-term expected emissions and utilities instead of just the current values, LEAD optimizes these metrics with a forward-looking perspective. The time complexity of LEAD depends on the size of the optimization problem per batch. While the problem size may grow with the number of drivers, practical techniques can effectively manage this complexity. For example, the optimization can be limited to drivers within a fixed range of the rider, as online assignment algorithms generally exclude distant drivers due to the high deadhead miles involved. This approach significantly reduces the problem size while maintaining solution quality.
@@ 4.2. Batched Emission-Aware Fair Assignment
2       Calculate $E_{v,r,b}$ using Equation (8);
3       Calculate $U_{v,r,b}$ using Equation (9);
@@ 11       Update $V_{T}(s_{t})$ using Equation (7);
5. Experimental Evaluation
In this section, we conduct comprehensive experiments to evaluate the performance of LEAD using several metrics, including emissions per trip, fairness in driver’s utility, and average wait times for riders. We compare LEAD against state-of-the-art algorithms across a range of values for key configuration paremeters of batch durations, the fairness parameter $\eta$, and the percentages of low-emission vehicles. The following outlines the key questions the evaluation addresses and summarizes our findings.
@@ 5. Experimental Evaluation
    How much emissions does LEAD reduce, especially compared to other baselines targeting emissions reduction?
    Outcome: LEAD reduces emissions by at least 56.6%, 23.5%, and 8.4%, compared to baseline algorithms that target improving fairness (LAF (Shi et al., 2021b)), reducing deadhead miles (CD), and minimizing carbon emissions (TORA  (Sahebdel et al., 2024)), respectively. 111 Section 5.1 describes the baseline algorithms in detail. This substantial decrease highlights LEAD’s effectiveness in minimizing environmental impact (see Section 5.2).
@@
    How much improvement in fairness does LEAD provide compared to the baseline targeting emissions reduction?
    Outcome: LEAD demonstrates a 150% improvement in fairness compared to TORA that balances minimizing system-wide emissions against the passenger wait times but does not explicitly target fairness across drivers (see Section 5.3).
@@
    How does LEAD’s rider wait time performance compare to the baseline algorithm that explicitly considers fairness?
@@
    Outcome: LEAD achieves 39.4% improvement over LAF while also ensuring that not degrading fairness for drivers by more than 28.0% (see Sec. 5.4).
5.1. Experimental Setup
Parameter settings. We evaluate the performance of LEAD and baseline algorithms across a range of experimental settings. Specifically, we evaluate the impact of parameter $\eta$ by varying its values between 0.1 gCO2/km and 10 gCO2/km. We also evaluate the impact of different batch durations by adjusting the batch duration values between 2 and 10 minutes. To define unique states, we divide the city into square tiles, each with a width of 1 km. The value functions for the reinforcement learning model are initialized to 0. We set the discount factor, $\gamma=0.9$, and the learning rate $\alpha=0.025$ in line with the evaluation settings used in previous research (Shi et al., 2021b), ensuring consistency in the comparative analysis. To introduce additional metrics for evaluating the system-level performance of LEAD compared to baseline algorithms, we classify vehicles emitting less than 135 gCO2/km as low-emission (LEVs) and those emitting more than 270 gCO2/km as high-emission (HEVs).
Ridesharing dataset. Our experiments use a publicly available dataset from RideAustin (RideAustin, 2017), a non-profit ridesharing service based in Austin, Texas. This dataset includes approximately 1.5 million trips over ten months in 2016 and 2017, collected in Austin. It contains detailed trip information such as pick-up and drop-off coordinates of each trip, vehicle make and model, and distances traveled before, during, and after each trip.
In our experiments, we use a subset of the RideAustin dataset, focusing on trips from December 1, 2016, to December 15, 2016. This subset includes 58,866 ridesharing trips served by 150 drivers. We augment this dataset by including carbon emission data for both deadhead miles and individual trips. Additionally, we include equity information for both drivers and riders, using E2-RideKit (Sahebdel et al., 2024), a publicly available toolkit designed for holistic evaluation of emissions in ridesharing platforms. We assume that riders have an expectation regarding the maximum time they are willing to wait before being assigned to a driver. Furthermore, drivers available within a 15-minute window are included in the set of available drivers for each batch. To assess the impact of replacing HEVs with LEVs, we created three different variants of original dataset. In each variant we randomly converted a certain percentage of the non-LEVs into electric vehicles (EVs). This process resulted in three datasets where the proportion of LEVs was increased to 10%, 15%, and 25%, respectively. We set the EV’s emission intensity to 63.35 gCO2 per kilometer. [^2] We used the Tesla Model Y as a representative EV, which has an energy efficiency rating of 16.9 kWh/100km. The unit distance emissions were calculated using the average carbon intensity of 408 gCO2eq/kWh for Austin, Texas.
Comparison algorithms. We evaluate the performance of LEAD by comparing it against a heuristic baseline algorithm as well as two state-of-the-art algorithms from prior works, which are outlined in detail below.
  • Closest Driver (CD): Within a given batch of requests, this algorithm sequentially matches ride requests in the batch to the nearest available driver. The matching process is conducted in a greedy manner, aiming to minimize the deadhead miles and reduce the rider’s wait time. Although this algorithm does not explicitly account for emissions during the assignment process, the focus on minimizing deadhead distances indirectly helps control ride emissions.
  • LAF (Shi et al., 2021b): This algorithm is designed to ensure a fair distribution of utility among drivers. The ride assignment process in LAF consists of four key stages: evaluating, assigning, guiding, and learning. Initially, the algorithm begins with evaluating a bipartite graph where edge weights represent trip prices between passengers and drivers. These weights are updated to reflect both current and future earnings based on historical trip prices observed in the area near the pick-up and the drop-off locations. During the assigning stage, LAF utilizes a bi-objective assignment algorithm. This algorithm aims to balance the aggregate utilities of drivers and the equity in their utilities. The learning stage involves refining weight updates through reinforcement learning. Finally, during the guiding stage, idle drivers will be guided to high-demand areas for their future assignments.
    Since LAF aims to maximize total utilities and ensure equity in utilities, it uses a single value function to track the expected achievable utilities in different city regions. For adaptation to the current work, we set the utility (or price) of a ride request $r$ served by driver $v$ as the difference between the trip and deadhead distances and evaluated weights in Algorithm 1 of (Shi et al., 2021b) accordingly. Additionally, during our experiments, the weight of trip utilities over different times of the day is the same.
  • TORA (Sahebdel et al., 2024): TORA is an online ride-assignment algorithm that aims to balance passenger waiting times with system-wide emissions. To the best of our knowledge, it is the only existing online ride assignment algorithm that explicitly considers emissions. TORA aims to reduce passenger waiting times by initially selecting the closest available driver. It then evaluates alternative available drivers by comparing their distances and the additional deadhead emissions they would generate relative to the closest driver. For each driver, TORA calculates an “Emission-to-Distance” (E2D) ratio, which is the ratio of the difference in deadhead emissions between a given driver and the closest driver to the difference in distances to the passenger between the two drivers. TORA selects the most suitable driver based on these E2D values. Despite the consideration of carbon emissions during the process of ride matching, TORA does not consider fairness in driver’s utilities.
Performance metrics. We evaluate the performance of LEAD and comparison algorithms using three primary performance metrics: per trip emissions, fairness in driver’s utility, and rider waiting time. These metrics provide a multifaceted view of how each algorithm performs across critical aspects of ridesharing services—environmental impact, driver equity, and the quality of experience of the riders. For fairness, we normalize our results to LAF algorithm’s outcomes when comparing different algorithms. However, in analyzing the effect of LEV penetration, we normalize the values to outcomes achieved by the LEAD algorithm without the addition of LEVs.
@@ 5.1. Experimental Setup
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of parameter $\eta$
@@
Figure 2. Emissions reduction performance: (a) emission per trip for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD for different batch durations and $\eta$ = 5, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$, for a batch duration of 5 minutes. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.2. Impact on Per-Trip Emissions
Key results. We first outline the key observations here and defer the insights and analysis of findings to the next part. As shown in Figure 2 (a), LEAD outperforms all the baseline algorithms across all different batch durations. The effectiveness of LEAD in reducing emissions per trip is evident, with reductions of at least 23.5% compared to the CD algorithm, 8.4% compared to the TORA algorithm, and 56.6% compared to the LAF algorithm. Notably, LEAD’s performance worsens as the batch duration increases. While the emissions per trip for CD and TORA slightly decrease with an increase in batch duration, LEAD achieves higher reductions at shorter batch duration, thereby increasing the margin of improvement. The maximum improvement of LEAD compared to TORA, CD, and LAF occurs at a batch duration of 2 minutes, with improvements of 14.6%, 31.4%, and 58.4%, respectively. Finally, LAF’s performance worsens with an increase in batch duration, e.g., emissions are 3% and 1% higher at 10 minutes batch duration than at 2 and 5 minutes batch duration.
As shown in Figure 2 (b), adding low-emission vehicles (LEVs) to the ridesharing fleet leads to a reduction in the average emissions per trip. Specifically, adding 10% LEVs results in a reduction of emissions by up to 21.8%. Increasing the proportion of LEVs to 15% reduces emissions by up to 32.5%. When the proportion of LEVs is increased to 25%, the emissions reduction achieves up to 37.9% demonstrating a significant improvement in environmental performance.
Figure 2 (c) shows that changing the fairness parameter $\eta$ has only a minor impact on the emissions. The maximum increase in emissions per trip is observed to be 0.8%, and this occurs when 10% of the vehicles are LEVs. As we increase $\eta$ to 10gCO2/km, the emissions only increase by 0.21% and 0.67% at 15% and 20% LEV percentages, respectively.
Analysis of findings. We next delve deeper into some results observed about LEAD’s impact on emissions.
  • Why does LEAD outperforms deadhead mile- and emission-aware algorithms CD and TORA?
    The primary reason that LEAD outperforms CD is that CD does not take into account the unit emissions of the vehicles when assigning requests to drivers. This omission means that CD prioritizes minimizing the distance between drivers and riders without factoring in the environmental impact of each vehicle’s emissions. When there is a wide variation in the emission rates across vehicles, the performance gap between CD and emission-aware algorithms like LEAD and TORA becomes more pronounced. Another reason that LEAD outperforms TORA and CD is its approach for the assignment of multiple requests within a batch. It assigns all the requests in each batch simultaneously, whereas TORA and CD assign the requests sequentially within each batch. This sequential approach limits their ability to optimize for emissions across the entire batch, as decisions are made one at a time rather than considering the batch as a whole.
  • Why does increasing the batch duration from 2 to 10 minutes increase emissions but improve fairness?
    The RideAustin dataset includes many rides with short trip distances. When the batch duration is short (e.g., 2 minutes), the system assigns requests more frequently allowing it to repeatedly assign rides to low-emission vehicles. This frequent assignment of low-emission vehicles reduces emissions but leads to unfairness since high-emission vehicles are underutilized. When the batch duration increases (e.g., to 10 minutes), the system processes more requests at once, however assigns only a subset of these to low-emission vehicles. The remaining requests are distributed among high-emission vehicles. This results in a higher average emission but fairer allocation of rides between low-emission and high-emission vehicles.
  • Why don’t the emissions per trip for LEAD increase as $\eta$ rises from $0.1$ to $10$ gCO2/km?
    The value of $\eta$ shows the grams of CO2 emitted for each kilometer increase in fairness. Within that range of values for $\eta$, LEAD carefully balances this tradeoff and only chooses the trips when a meaningful improvement in fairness can be achieved.
Key takeaways. LEAD achieves up to 14.6% emissions reduction compared to the state-of-the-art emission-aware algorithm TORA. While additional reductions may be possible by optimizing parameters such as batch duration and $\eta$, it will introduce tradeoffs with our other objectives of increasing driver’s fairness and reducing rider’s wait time.
5.3. Impact on Driver’s Fairness
Key results. As shown in Figure 3 (a), no other algorithm, including LEAD, beats LAF in terms of fairness performance. This happens since LAF only prioretize fairness in its assignments. The next best fairness performance is shown by LEAD, as it reaches 71.9% of the fairness achieved by LAF at the batch duration of 10 minutes. Meanwhile, TORA is the worst performing algorithm as it achieves between 19.1% to 39.0% of LAF’s fairness performance at batch durations of 2 minutes and 10 minutes, respectively.
As shown in Figure 3 (b), incorporating 25% LEVs has the most significant impact on fairness. Specifically, with 25% LEVs, fairness improves by 40% for batch durations of 2 minutes compared to the baseline ridesharing fleet with LEAD. In contrast, increasing the proportion of LEV to 10% or 15% results in a decrease in fairness over all batch durations.
Figure 3 (c) shows the effect of $\eta$ on fairness, providing us with two key observations. First, increasing the value of $\eta$ increases fairness. For example, at an $\eta$ value of 0.1 gCO2/km when 25% of the fleet consists of LEVs, fairness increases by 21%, while a much higher $\eta$ value of 10 gCO2/km leads to an 45% increase in fairness. Second, fairness decreases at lower LEV percentages but improves significantly at higher percentages. For example, fairness improved by at least 60% and up to 70% when the percentage of LEVs was increased from 15% to 25%.